### v2

# Speedups of linearly recurrent subshifts

Abstract
A speedup, like a time change in discrete time dynamics, is a way of moving faster through the orbits of a dynamical system. Linearly recurrence is a stronger form of minimality for subshifts, shared by e.g. all primitive substitution shifts and Sturmian shifts associated with rotation numbers of bounded type. We prove that the homeomorphic speedup of a linearly recurrent two-sided subshift is again linearly recurrent.
Mathematics Subject Classification (2010): Primary: 37B10, Secondary: 37B05
Keywords. speedup, linear recurrence, subshift, word-complexity.
1 Introduction
A speedup of a discrete time dynamical system $(X,T)$ is another dynamical system that emerges by going faster through the $T$-orbits. Namely, given a jump-function $p:X\to{\mathbb{N}}=\{1,2,3,\dots\}$, and the speedup is $S:X\to X$, $S(x)=T^{p(x)}(x)$. Speedups can be viewed as a kind of discrete time-changes for continuous time flows, cf. [12]. The jump-function is a weaker form of the orbit cocycle of orbit equivalence, see e.g. [7, page 23]. Speedups were considered from a measure-theoretic point of view in [3] (its relation to cutting-and-stacking) and [11] (dimension groups for speedups). More topological approaches were given in [1, 4, 6].
Throughout this paper, we assume that $X$ is a Cantor set and $T:X\to X$ is continuous and invertible. We assume that the jump-function is continuous, injective (i.e., $T^{p(x)}(x)\neq T^{p(y)}(y)$ whenever $x\neq y$) and is such that $S$ is again invertible. Since the space $X$ is compact, this implies in particular that $p$ is bounded, so that
$p_{\max}:=\max\{p(x):x\in X\}\ \text{ exists,}$ (1)
and that every $T$-orbit falls apart in a finite number of $S$-orbits. The maximum $c$ of these numbers of $S$-orbits is called the orbit number and $c\leq\max_{x}p(x)$. The result [1, Proposition 2.2] (and also [4, Proposition 4.1.2]) tells us that if $(X,T)$ is a minimal Cantor system with a continuous speedup $S=T^{p}$, then $p$ is continuous if and only if it is bounded.
A general question is which properties the speedup $(X,S)$ inherits from $(X,T)$. Transitivity and minimality are not inherited in general. To illustrate this, assume that $\{X_{1},X_{2}\}$ is a non-trivial clopen partition of the Cantor set $X$. If $p(x)$ is the first return time to the partition element that $x$ belongs to, then $S=T^{p}$ has $X_{1}$ and $X_{2}$ as invariant subsets. For this reason, we will assume throughout that our speedups are transitive. However, the transitive speedup of a minimal system is again minimal, see Proposition 2.1.
In the context of symbolic systems (i.e., subshifts and odometers), it is easy to show that speedups of subshifts of finite type (SFTs) are again SFTs, and speedups of sofic shifts are sofic, see Proposition 2.2. Ash et al. [1, 6] showed, among other results, that minimal speedups of substitution shifts are substitution shifts, minimal speedup of odometers are odometers. However, speedups of Toeplitz shifts need not be Toeplitz shifts, see [2].
Let $(X,\sigma)$ be a subshift, where $\sigma$ denotes the left-shift. We say that $(X,\sigma)$ is linearly recurrent if there is $L\in{\mathbb{N}}$ such that for every $x\in X$ each every finite word $w$ appearing in $X$ appears in $x$ with gap $\leq L|w|$ shifts. The aim of this paper is to prove the following theorem.
Theorem 1.1
Let $S=\sigma^{p}$ be the homeomorphic transitive speedup of a two-sided subshift $(X,\sigma)$. Then $\sigma$ is linearly recurrent if and only if $S$ is linearly recurrent.
Only for two-sided shifts, a speedup can be homeomorphic, but by passing from a one-sided to the two-sided shift with the same language, one can easily extend Theorem 1.1 to one-sided shifts.
The next section gives the relevant definitions, some related results, and estimates the word-complexity of speedups. Section 3 investigates the structure of non-abelian finite group-extensions and gives results on essential values and minimality needed for the main theorem. Finally, in Section 4 we prove the main theorem.
2 Preliminaries and miscellaneous results
As mentioned in the introduction, speedups of transitive (Cantor) systems need not be transitive, but we have the following:
Proposition 2.1
The continuous homeomorphic transitive speedup of a minimal system is minimal.
[…]
2.1 Subshifts of finite type and sofic shifts
Let $\mathcal{A}$} be a finite alphabet, and equip the left-shift on the space $\mathcal{A}$$^{{\mathbb{Z}}}$ of two-sided infinite sequences with product topology. Let $\sigma:{\mathcal{A}}^{\mathbb{Z}}\to{\mathcal{A}}^{\mathbb{Z}}$, $\sigma(x)_{i}=x_{i+1}$, be the the left-shift. We say that $(X,\sigma)$ is a subshift if $X$ is $\sigma$-invariant (in fact, $\sigma(X)=X$) closed non-empty subset of $\mathcal{A}$$^{{\mathbb{Z}}}$. Let ${\mathcal{L}}(X):=\{w\in{\mathcal{A}}^{n}:n\in{\mathbb{N}},w\text{ appears as a subword in some }x\in X\}$ be the language of the subshift. Formally, also the empty word $\epsilon$ belongs to ${\mathcal{L}}(X)$, but we don’t need it, and including it only makes formulas more cumbersome.
A subshift of finite type (SFT) $(X,\sigma)$ on a finite alphabet $\mathcal{A}$} is characterized a vertex-labeled transition graph $G=(V,E)$, where the vertex set $V={\mathcal{A}}$ and the directed edge set $E$ is unlabeled. Infinite word paths in this graph correspond bijectively to the elements $x\in X$, via the labeling of the vertices.
A sofic shift $(X,\sigma)$ on a finite alphabet $\mathcal{A}$} is characterized by an edged-labeled transition graph $G=(V,E)$. The vertex set is unlabeled but each (directed) $e\in E$ gets a label from $\mathcal{A}$}, where (contrary to SFTs) the same label can be assigned to multiple edges. Infinite word paths in this graph correspond bijectively to the elements $x\in X$, via the labeling of the edges.
Proposition 2.2
The injective continuous speedup of a subshift of finite type is a subshift of finite type. The injective continuous speedup of a sofic shift is sofic.
Proof. Let $N$ be minimal such that $p$ is constant on every centered $2N+1$-cylinder. Set $M=2\max\{p_{\max},N\}+1$ and consider the alphabet $\tilde{\mathcal{A}}={\mathcal{A}}^{M}$. The SFT $(X,\sigma)$ over $\mathcal{A}$} can be represented as an $M$-block-code shift over $\tilde{\mathcal{A}}$, with its vertex-labeled transition graph $G=(V=\tilde{\mathcal{A}},E)$. Speeding up $\sigma$ results into a transition graph $G^{\prime}=(V^{\prime},E^{\prime})$ where $V^{\prime}=V=\tilde{\mathcal{A}}$ and for each $v\in V$, the edges $e\in E^{\prime}$ outgoing from $v$ are the paths of length $p(v)$ starting at $v$. That is, $(X,S=\sigma^{p})$ has a vertex-labeled transition graph, so it is an SFT.
The proof for sofic shifts is the same, only with edge-labeled transition graphs. $\Box$
2.2 Word-complexity of speedups
In the above notation of subshifts, linear recurrence means that for every $w\in{\mathcal{L}}(X)$ and $x\in X$ and $i\in{\mathbb{Z}}$, there is $k\in\{0,1,\dots,L|w|\}$ such that $x_{i+k+1}\dots x_{i+k+|w|}=w$. Linear recurrence is preserved by topological conjugacy, although the constant $L$ can change. Linearly recurrent shifts are automatically minimal and in fact uniquely ergodic.
Let $\mathfrak{p}_{\sigma}(n)=\#\{w\in{\mathcal{L}}(X):|w|=n\}$ be the word-complexity of $(X,\sigma)$. Linearly recurrent shifts have sublinear complexity: $\mathfrak{p}_{\sigma}(n)\leq Ln$. If $(X,\sigma)$ is linearly recurrent with constant $L$, then its word-coplexity $\mathfrak{p}_{\sigma}(n)\leq Ln$ as well, but there are many subshifts of linear word-complexity that are not linearly recurrent, e.g. all Sturmian subshifts associated to a rotation number of unbounded type. So the following proposition is, for linear recurrent shifts, a direct consequence of Theorem 1.1, but it works much more generally and has a much easier proof than Theorem 1.1.
Proposition 2.3
Given a homeomorphic speedup $S=\sigma^{p}$ of a subshift $(X,\sigma)$, there is a constant $K$ such that the word-complexity of $S$ satisfies $\mathfrak{p}_{S}(n)\leq K\,\mathfrak{p}_{\sigma}(p_{\max}\,n)$ for $p_{\max}=\max\{p(x):x\in X\}$ from (1).
Proof. Since the jump-function is continuous and the space $X$ is compact, there is $N\in{\mathbb{N}}$ such that $p$ is constant on each $N$-cylinder.
Let $Z=[z_{1}\dots z_{n}]\subset X$ be any $n$-cylinder. If a sped-up orbit visits position $k\in[N+1,n-N]$ of this cylinder set, then the other positions in [N+1,nN][N+1,n-N] are fixed as well. We make no such uniqueness claim for positions in $[1,N]$ and [nN+1,n][n-N+1,n], but these are at most ${\mathcal{A}}^{2N}$ patterns of positions. Hence, associated to $Z$, there are at most $p_{\max}\#{\mathcal{A}}^{2N}$ $n/p_{\max}$-cylinders. This gives $\mathfrak{p}_{S}(n)\leq K\,\mathfrak{p}_{\sigma}(p_{\max}n)$ for $K=\#{\mathcal{A}}^{2N}p_{\max}$ where $p_{\max}$ is from (1). $\Box$
It follows that if $(X,\sigma)$ has linear, quadratic, cubic, etc., or even exponential word-complexity, so does the speedup. For positive entropy shifts, this result just reproduces Abramov’s formula $h_{top}(Y,S)=(\int_{X}p\,d\mu)h_{top}(X,\sigma)$ as in [13] and [1, Proposition 2.9].
3 Finite non-abelian group extensions
Let $T:X\to X$ be a minimal homeomorphism on a Cantor set. For a finite [^1] Much of this section also holds for compact groups, but as we only apply it to a group of permutations in Section 4, we restrict ourselves to finite groups. group $G$, consider a skew-product
$F:X\times G\to X\times G,\qquad(x,g)\mapsto(T(x),g\cdot\phi(x))$
for a continuous skew-function $\phi:X\to G$. Write
[…]
where the final -1 in the third line refers to the inverse in the group.
Define for each $x\in X$ the set
$G_{x}:=\{g\in G:\exists n_{k}\nearrow\infty\text{ such that }\phi^{n_{k}}(x)\equiv g\text{ and }T^{n_{k}}(x)\to x\}.$ (3)
If $g,g^{\prime}\in G_{x}$, say with corresponding sequence $(n_{k})_{k\in{\mathbb{N}}}$ and $(n^{\prime}_{l})_{l\in{\mathbb{N}}}$, then, by continuity of $\phi^{n_{k}}$,
$\phi^{n_{k}+n^{\prime}_{l}}(x)=g^{\prime}\cdot\phi^{n_{k}}(T^{n^{\prime}_{l}}(x))\to g^{\prime}\cdot g,\quad\text{ as }l\to\infty.$ (4)
This holds for every $k$. Hence, by continuity of each $T^{n_{k}}$, there is a sequence $k_{l}\nearrow\infty$ such that $T^{n_{k_{l}}+n^{\prime}_{l}}(x)\to x$. This shows that $G_{x}$ is closed under composition.
Remark 3.1
One can just as well argue that, by continuity of $\phi^{n^{\prime}_{l}}$,
$\phi^{n_{k}+n^{\prime}_{l}}(x)=\phi^{n^{\prime}_{l}+n_{k}}(x)=g\cdot\phi^{n^{\prime}_{l}}(T^{n_{k}}(x))\to g\cdot g^{\prime},\quad\text{ as }l\to\infty.$
However, this still doesn’t imply that $G_{x}$ is abelian, because $(n_{k}+n^{\prime}_{l})_{l\in{\mathbb{N}},k\in K}$ and $(n^{\prime}_{l}+n_{k})_{k\in{\mathbb{N}},l\in L}$ might be entirely different sequences for infinite sets $K,L\subset{\mathbb{N}}$, with only finite overlap. Letting $k\to\infty$ and $l\to\infty$ at different speeds can still lead to different limit points in $G$. It seems that if $T$ is equicontinuous, then $G_{x}$ is an abelian subgroup.
Because $G$ is a finite group, every $g\in G$ has finite order, say order $d$, so $g^{d}=e$ is the identity. Applying (4) d2d-2 and d1d-1 times with $g^{\prime}=g$, we obtain that $g^{d-1}=g^{-1}\in G_{x}$ and $g^{d}=e\in G_{x}$. This proves for each $x\in X$ that $G_{x}$ is a subgroup of $G$.
Lemma 3.1
Let $x,y\in X$ be two points in the same $T$-orbit. Then $G_{x}$ and $G_{y}$ are conjugate subgroups of $G$, i.e., there is $g$ such that $G_{y}=g\cdot G_{x}\cdot g^{-1}$. We can in fact take $g=\phi^{n}(y)$ where $n\in{\mathbb{Z}}$ is such that $T^{n}(y)=x$.
Proof. Take $n\in{\mathbb{Z}}$ such that $T^{n_{k}}(y)\to x$ and $g=\phi^{n}(y)$. Now for $h\in G_{x}$ arbitrary, take $n^{\prime}_{l}\nearrow\infty$ such that $T^{n^{\prime}_{l}}(x)\to x$ and $\phi^{n^{\prime}_{l}}(z)\equiv h$. Then
$\phi^{-n+n^{\prime}_{l}+n}(y)$ $=$ $\phi^{n}(y)\cdot\phi^{n^{\prime}_{l}}(x)\cdot\phi^{-n}(T^{n^{\prime}_{l}}(x))=g\cdot h\cdot\phi^{-n}(T^{n^{\prime}_{l}}(x))$
$\to_{l\to\infty}$ $g\cdot h\cdot\phi^{-n}(x)=g\cdot h\cdot g^{-1},$
because $\phi^{-n}(x)=\phi^{-n}(T^{n}y)=(\phi^{n}(y))^{-1}=g^{-1}$ by (2). Since this holds for all $h\in G_{x}$, we have $G_{y}\supset g\cdot G_{x}\cdot g^{-1}$, and in particular, $G_{y}\geq\#G_{x}$. Reversing the roles of $x$ and $y$ gives the required equality. $\Box$
Proposition 3.1
Assume that $T:X\to X$ is minimal. The map $x\mapsto G_{x}$ is piecewise constant.
Proof. Let $x\in X$ be “minimal” in the sense that for no $y\in X$, $G_{y}$ is a proper subgroup of $H:=G_{x}$. For each $g\in G$, we can try to find a sequence $(r^{g}_{k})_{k\in{\mathbb{N}}}$ such that $\phi^{r^{g}_{k}}(x)\equiv g$ and $d(T^{r^{g}_{k+1}}(x),x)\leq\frac{1}{2}d(T^{r^{g}_{k}}(x),x)$ for all $k$. If $g\notin H$, then such a sequence is necessarily finite, so there is a largest integer $N_{g}$ in this sequence. (If $\phi^{n}(x)\neq g$ for all $n\geq 0$, then we set $N_{g}=0$.)
Let $N=\max\{N_{g}:g\notin H\}$ and let $U_{x}$ be a clopen neighborhood of $x$ (in fact, a cylinder set if $(X,T)$ is a subshift), such that $T^{n}(x)\notin U_{x}$ for all $1\leq n\leq N$. It follows that if $T^{n}(x)\in U_{x}$, then $\phi^{n}(x)\in H$. Let $V\subset U_{x}$ be any maximal clopen set such that its return time $r_{V}$ to $U_{x}$ is constant and $\phi^{r_{V}}|_{V}\equiv g_{V}$ is constant, as well. By minimality, there is $n\in{\mathbb{N}}$ such that $T^{n}(x)\in V$, so $g:=\phi^{n}(x)\in H$. But then $T^{n+r_{V}}(x)\in U_{x}$, so that also $\phi^{n+r_{V}}(x)=g\cdot g_{V}\in H$, so $g_{V}\in H$.
Since $g_{V}\in H$ for every return domain $V$ of $U_{x}$, it follows $\phi^{n}(y)\in H$ whenever $y,T^{n}(y)\in U_{x}$, and therefore $G_{y}\subset G_{x}=H$. But the choice of $x$ at the beginning of the proof means that $G_{y}=G_{x}$ for all $y\in U_{x}$. Therefore $y\in X_{H}$ for all $y\in U_{x}$, so $x$ is an interior point of $X_{H}$, and because $x\in X_{H}$ was arbitrary, $X_{H}$ is open.
By minimality, there is $\ell\in{\mathbb{N}}$ such that $X=\bigcup_{i=0}^{\ell-1}U_{x}$, and by Lemma 3.1, $G_{z}=g\cdot H\cdot g^{-1}$ for each $z\in T^{i}(U_{x}\cap A_{i})$, $i=0,\dots,\ell-1$, where $A_{i}$ is any set on which $\phi^{i}\equiv g$ is constant. This proves that $x\mapsto G_{x}$ is piecewise constant. $\Box$
Corollary 3.1
Let $T:X\to X$ be a minimal Cantor system, and $x,y\in X$ arbitrary. Then $G_{x}$ and $G_{y}$ are conjugate subgroups of $G$, i.e., there is $g$ such that $G_{y}=g\cdot G_{y}\cdot g^{-1}$.
Proof. This follows immediately from Lemma 3.1 and the continuity of $x\mapsto G_{x}$. In detail: by minimality, we can choose $n_{k}\nearrow\infty$ such that $T^{n_{k}}(y)\to x$ and since $G$ is finite, we can assume that $\phi^{n_{k}}(y)\equiv g$. Now for $h\in G_{x}$ arbitrary, take $k$ so large, i.e., $z:=T^{n_{k}}(y)$ so close to $x$ that $h\in G_{z}$. Next take $n^{\prime}_{l}\nearrow\infty$ such that $T^{n^{\prime}_{l}}(z)\to z$ and $\phi^{n^{\prime}_{l}}(z)\equiv h$. Then
$\phi^{-n_{k^{\prime}}+n^{\prime}_{l}+n_{k}}(y)$ $=$ $\phi^{n_{k}}(y)\cdot\phi^{n^{\prime}_{l}}(z)\cdot\phi^{-n_{k}}(T^{n^{\prime}_{l}}(z))=g\cdot h\cdot\phi^{-n_{k}}(T^{n^{\prime}_{l}}(z))$
$\to_{l\to\infty}$ $g\cdot h\cdot\phi^{-n_{k}}(z)=g\cdot h\cdot g^{-1},$
because $\phi^{-n_{k}}(z)=\phi^{-n_{k}}(T^{n_{k}}y)=(\phi^{n_{k}}(y))^{-1}=g^{-1}$ by (2). Since this holds for all $h\in G_{x}$, we have $G_{y}\supset g\cdot G_{x}\cdot g^{-1}$, and in particular, $G_{y}\geq\#G_{x}$. Reversing the roles of $x$ and $y$ gives the required equality. $\Box$
The sets $G_{x}$ play the role of local essential values. We say that $h\in G$ is an essential value if for every $A\subset X$ with $\mu(A)>0$ there is $n\in{\mathbb{Z}}$ such that
$\mu\left(A\cap T^{-n}A\cap\{x\in X:\phi^{n}(x)=h\}\cap\{x\in X:T^{n}(x)\neq x\})\right)>0.$
For abelian groups $G$, $G_{x}$ coincides with the set of essential values, but for non-abelian groups, the essential values may not be more than the centralizer of $G$, or the intersection $\bigcap_{x\in X}G_{x}$, and thus $G$ is of less use.
The collection of essential values forms a subgroup of $G$, see [14].
Lemma 3.2
Let $T:X\to X$ be minimal with an ergodic $T$-invariant measure $\mu$. Let $F(x,g)=(T(x),g\cdot\phi(x))$ be a continuous skew-product, for a finite group $G$. Then there is $x\in X$, a neighbourhood $U$ of $x$ and an ergodic $F$-invariant measure $\rho$ such that $\rho|_{U\times H}$ is a multiple of $\mu|_{U}\otimes\nu_{H}$, where $H=G_{x}$ and $\nu_{H}$ is right Haar measure (i.e., counting measure) on $H$.
Proof. The proof of Proposition 3.1 implies that there is $x\in X$ and a neighbourhood $U=U_{x}$ of $x$ such that if $T^{n}(y)\in U$ for some $y\in U$, then $\phi^{n}(y)\in H$. Let $T_{U}=T^{\tau}$ be the first return map to $U$, where $\tau(y)=\min\{n\geq 1:T^{n}(y)\in U\}$ is the first return time of $y\in U$ to $U$. Then $\mu_{U}:=\frac{1}{\mu(U)}\mu|_{U}$ is $T_{U}$-invariant and ergodic because $\mu$ is $T$-invariant and ergodic. Let
$F_{U}:U\times H\to U\times H,\qquad(y\,,\,g)\mapsto(T_{U}(y)\ ,\ g\cdot\phi^{\tau(y)}(y))$
be the corresponding skew-product; it preserves $\mu_{U}\otimes\nu_{H}$.
To show ergodicity, assume that $Y\subset U\times H$ is invariant under the skew-product and $\mu_{U}\otimes\nu_{H}(Y)>0$. Since $\mu_{U}$ is ergodic, the projection of $Y$ onto $U$ has full $\mu_{U}$-measure.
Assume now by contradiction that $\mu_{U}\otimes\nu_{H}(Y^{c})>0$. Both $Y$ and $Y^{c}=(U\times H)\setminus Y$ are invariant, so their projections are $T_{U}$-invariant sets of $U$ of positive measure, and by ergodicity of $\mu$, these projections are $U\bmod\mu_{U}$. Therefore there is a set $A\subset U$ and $h_{1},h_{2}\in H$ such that
$\mu\otimes\nu_{H}(Y\cap(A\times\{h_{1}\}))>0\quad\text{ and }\quad\mu\otimes\nu_{H}(Y\cap(A\times\{h_{2}\}))=0.$ (5)
(We could write $T$ here instead of $T_{U}$ because $T^{n}(y)\in U$ for $y\in U$ is equivalent to $T^{n}(y)=T^{n^{\prime}}_{U}(y)$ for some $|n^{\prime}|\leq|n|$.) Take $h=h_{1}^{-1}\cdot h_{2}\in H$, so by definition of essential value, there is $n\in{\mathbb{Z}}$ such that $\mu(B)>0$ for $B:=A\cap T^{-n}(A)\cap\{x\in X:\phi^{n}(x)=h,x\neq T^{n}(x)\}$. But then
$\mu\otimes\nu_{H}(\{(T^{n}(x),h_{1}\cdot\phi^{n}(x)):x\in B\})>0,$
but
$\{(T^{n}(x)\,,\,h_{1}\cdot\phi^{n}(x)):x\in B\}\subset A\times\{h_{2}\},$
contradicting the choice of $A$ and $h_{2}$ in (5). Hence $\mu_{U}\otimes\nu_{H}$ is ergodic.
Finally, using Kac’ formula, we spread this mass over the whole skew-product space to obtain the required $F$-invariant measure $\rho$:
$\rho(\,\cdot\,):=\frac{1}{\Lambda}\sum_{n\geq 1}\mu_{U}\otimes\nu_{H}\left(F^{-n}(\,\cdot\,)\cap\{(y,g)\in U\times H:n<\tau(y)\}\right),$
for normalising constant $\Lambda:=\sum_{n\geq 1}n\mu_{U}(\{y:\tau(y)=n\})$. This formula gives $\rho|_{U\times H}=\frac{1}{\Lambda}\mu_{U}\otimes\nu_{H}$, and the ergodicity of $\rho$ follows from the ergodicity of $\mu_{U}\otimes\nu_{H}$. $\Box$
The following result is in essence due to Furstenberg [10, Section 2 and Theorem 4.1], who proved it for toral automorphisms. The idea was used to prove the unique ergodicity (and minimality) of arbitrary group extensions over uniquely ergodic bases.
Proposition 3.2
Let $(X,T,\mu)$ be uniquely ergodic and $G$ a compact group with (right) Haar measure $\nu_{G}$. Let the group extension $S:Y\to Y$ defined on $Y:=X\times G$ as $F(x,g)=(T(x),g\phi(x))$ for some $\phi:X\to G$. If $F$ is ergodic w.r.t. $\nu=\mu\otimes\nu_{G}$, then $F$ is uniquely ergodic.
Proof. Let $(x,g)\in Y$ be the $\mu\otimes\nu_{G}$-typical point, so it satisfies Birkhoff’s Ergodic Theorem w.r.t. every continuous function $\varphi:Y\to{\mathbb{R}}$. For any $h\in G$, $\varphi_{h}$ defined by $\varphi_{h}(g)=\varphi(hg)$ is continuous too, so $(x,hg)$ is $\nu$-typical w.r.t. $\varphi$ because $(x,g)$ is $\nu$-typical w.r.t. $\varphi_{h}$. It follows that there is a subset $W\subset X$ with $\mu(W)=1$ such that $W\times G$ consists entirely of $\nu$-typical points.
If $\nu^{\prime}$ was another ergodic $S$-invariant probability measure, then the argument above gives a set $W^{\prime}\subset X\setminus W$ with $\nu^{\prime}(W)=1$ such that $W^{\prime}\times G$ consist entirely of $\nu^{\prime}$-typical points. Then the projected measure $\mu^{\prime}$ on $X$ defined by $\mu^{\prime}(A)=\nu^{\prime}(A\times G)$ is $T$-invariant, and satisfies $\mu^{\prime}(W^{\prime})=1$. But $W$ and $W^{\prime}$ are disjoint, so $\mu\neq\mu^{\prime}$, contradicting that $T$ is uniquely ergodic. $\Box$
4 Linear recurrence of speedups
A major tool in the proof of our main result Theorem 1.1 are return words. A word $R$ is called a return word of $w$ if $Rw$ begins and ends with $w$, but there are no other occurrences of $w$ in $Rw$. Linear recurrence implies that each return word $R$ has length $|R|\leq L|w|$ and there are at most $L(L+1)^{2}$ return words, see [9] and e.g. [7, Theorem 4.4]. Let us denote the return words of $w$ by $R_{1},\dots,R_{N}$ for $N\leq L(L+1)^{2}$. Furthermore, each return word reoccurs in $x\in X$ with bounded gap, and also this gap depends only on $L$. That is, there is $L^{\prime}$ independent of $w$ such that
$\max_{i\in{\mathbb{Z}}}\min\{n\geq 1:R_{e_{i+n}}=R_{e_{i}}\}\leq L^{\prime},$
where ${\mathcal{R}}:=\dots R_{e_{-2}}R_{e_{-1}}R_{e_{0}}R_{e_{1}}R_{e_{2}}\dots\in X_{w}$ is the concatenation of return words representing $x\in X$. In fact, $(X_{w},\sigma)$, the so-called derived shift of $(X,\sigma)$, is a linearly recurrent shift by itself. Despite there being infinitely many words $w$, there are only finitely many different derived shifts up to conjugacy, see [8].
Remark 4.1
An equivalent approach equivalent to return words is by means of cutting-and-stacking. The return words $R_{i}$ are then stacks and the letters in $\mathcal{A}$} that $R_{i}$ are composed of can be read off on the stack, bottom to top.
An outline of the proof of why linear recurrence of $(X,\sigma)$ implies linear recurrence of the speedup starts as follows. Each bi-infinite sequence $x\in X$ can be written as concatenation ${\mathcal{R}}=\dots R_{e_{-2}}R_{e_{-1}}R_{e_{0}}R_{e_{1}}R_{e_{2}}\dots$ of return words associated to some subword $w$ of $x$. We take $w$ so long that the return words contain the essential information of the $S$-orbit that pass through it. That is, the $c$ different $S$-orbits that $\operatorname{orb}(x)$ splits into, divides the entries of $R_{i}$ into $c$ subsets, characterized by so-called entry positions. For each pair of concatenated return words $R_{e_{i}}R_{e_{i+1}}$, the transit of the $j$-th $S$-orbit, starting at the $j$-th entry position in $R_{e_{i}}$, uniquely selects an entry position in $R_{e_{i+1}}$, say $j^{\prime}$. Doing this for all $j\in\{1,\dots,c\}$, we get a permutation $\psi_{R_{e_{i}}R_{e_{i+1}}}:\{1,\dots,c\}\to\{1,\dots,c\}$. This leads to a group extension, i.e, skew-product of $X_{w}\times{\mathfrak{S}}$, where $X_{w}$ is the shift space expressed in return words of $w$, and $\mathfrak{S}$} is the group of permutations of $\mathfrak{S}$}. This is the first step of the proof, which we formalize in the following construction.
Construction of the group extension: Let $(X,\sigma)$ be a two-sided shift, and let $S=\sigma^{p}$ be the speedup, where the jump-function $p:X\to{\mathbb{N}}$ is continuous. Since $X$ is compact, there is $K\in{\mathbb{N}}$ so that $p$ is constant on cylinders with anticipation and history $K$ (i.e., centered $2K+1$-cylinders). Consider $x\in X$ and a word $w$ in $x$ of length $|w|\geq p_{\max}+4K+2$.
Let $c\leq p_{\max}$ be the orbit number of $S$, so there are $c$ different $S$-orbits passing through $w$. Inside $w$, we can identify $c$ entry positions $j$ (where $2K+1<j<2K+p_{\max}$, for $p_{\max}$ as in (1)), that are the entry points of the $c$ distinct $S$-orbits inside $w_{2K+2}\cdots w_{2K+1+p_{\max}}$.
Let
$\hat{w}=w_{2K+1}\dots w_{|w|-2K}$
be the central subword of length $|\hat{w}|=|w|-4K$. If the initial position of an $S$-orbit in $\hat{w}$ is known, then the whole $S$-orbit inside $\hat{w}$ is fixed.
Furthermore, each concatenated pair $RR^{\prime}$ of return words induces a permutation
$\psi_{RR^{\prime}}:\{1,\dots,c\}\to\{1,\dots,c\}$
by setting $\psi_{RR^{\prime}}(s)=j^{\prime}$ if the $S$-orbit passing through the $j$-th entry position in the prefix word $w$ of $Rw$ also passes through the $j^{\prime}$-th entry position of the prefix word $w$ in $R^{\prime}w$. Let $\mathfrak{S}$} denote the collection of permutations of $\{1,\dots,c\}$. Thus, if the $c$ orbits (numbered $1,\dots,c$) enter the prefix word $w$ of the “first” letter $R_{e_{0}}w$ of $\mathcal{R}$} from the left to the right (which is indicated as the indentity permutation $e\in{\mathfrak{S}}$), then they enter the prefix word $w$ of $R_{e_{1}}w$ as $\psi_{R_{e_{0}}R_{e_{1}}}$, and they enter the prefix word $w$ of $R_{e_{1}}w$ as $\psi_{R_{e_{0}}R_{e_{1}}}\psi_{R_{e_{1}}R_{e_{2}}}$, etc. Thus the entries of the distinct orbits is described by a skew-product
$F:X_{w}\times{\mathfrak{S}}\to X_{w}\times{\mathfrak{S}},\qquad({\mathcal{R}},s)\mapsto(\sigma({\mathcal{R}}),\ s\cdot\psi_{R_{e_{0}}R_{e_{1}}}).$
Example: We illustrate the construction of the group extension by an example of the Fibonacci substitution shift. The Fibonacci substitution shift is the shift-orbit closure of the fixed point $\rho$ of the Fibonacci substitution $\chi:0\to 01,\ 1\to 0$:
$\rho=0\ \underbrace{10\ 01\ 0}_{R^{\prime}}\underbrace{10\ 0}_{R}\underbrace{1001\ 0}_{R^{\prime}}\underbrace{10010}_{R^{\prime}}\underbrace{10\ 0}_{R}\underbrace{10010}_{R^{\prime}}\underbrace{100}_{R}\underbrace{1001\ }_{R^{\prime}}\dots$ (6)
This is a one-sided sequence, and $\sigma^{-1}(\rho)$ has two preimages in $\overline{\operatorname{orb}_{\sigma}(\rho)}$, but this issue disappears when looking at the two-sided version.
Let $w=1001$, then there are two return words $R=100$ and $R^{\prime}=10010$, see (6). Assume the jump-function is $p\equiv 2$. Then $|w|$ is too short to precisely implement the above (in fact, one return word is even shorter than $w$ itself), but we can nonetheless declare $1$ and $2$ as the entry positions. Since $|R|$ and $|R^{\prime}|$ are both odd, $\psi_{RR^{\prime}}=\psi_{R^{\prime}R}=\psi_{R^{\prime}R^{\prime}}=(12)$ always flips these positions.
If $p\equiv 3$, then $\psi_{RR^{\prime}}=e$, the identity, because $|R|=3$, but $\psi_{R^{\prime}R}=\psi_{R^{\prime}R^{\prime}}=(123)$.
To continue the preparation for the proof of the main theorem, we state Lemma 4.1, which basically says that the skew-product we constructed is linearly recurrent.
Lemma 4.1
There is $L^{*}$ depending only on $L$ such that for every $w\in{\mathcal{L}}(X)$ and ${\mathcal{R}}\in X_{w}$, every $s\in{\mathfrak{S}}$ reoccurs within $L^{*}|w|$ iterates of $F$, i.e., there is $1\leq n\leq L^{*}|w|$ such that $F^{n}({\mathcal{R}},s)=(\sigma^{n}({\mathcal{R}}),s)$.
Before proving Lemma 4.1 formally, we give an outline of the proof. First, we study the shapes of the transitive components of $X_{w}\times{\mathfrak{S}}$. To find such component, choose $(\hat{x},\hat{s})\in X_{w}\times{\mathfrak{S}}$ and take its orbit closure, say $Y$. By minimality of $(X_{w},\sigma)$, $Y$ projects to the whole of $X_{w}$ in the base direction, but in the fiber direction, for each $x\in X_{w}$, $Y$ projects to $\{x\}\times s_{x}{\mathfrak{S}}_{x}$, where ${\mathfrak{S}}_{x}$ is a subgroup of $\mathfrak{S}$}, the analog of $G_{x}$ in (3) and $s_{x}\in{\mathfrak{S}}$. The assignment $x\mapsto{\mathfrak{S}}_{x}$ is continuous, and piecewise constant because $\mathfrak{S}$} is discrete. The proof of Proposition 3.1, in fact, gives a (cylinder) subset $U\subset X_{w}$, such that both $s_{x}\equiv e$ and ${\mathfrak{S}}^{\prime}:={\mathfrak{S}}_{x}$ is constant for all $x\in U$. This means that the first return map to $U$ is in fact the derived shift $(X_{w^{\prime}},\sigma)$.
Replacing the shift $\sigma$ on $X_{w}$ by the first return map to $U$, we can assure that the skew-product on $(X_{w^{\prime}},{\mathfrak{S}}^{\prime})$ is transitive and ergodic by Lemma 3.2. Now Proposition 3.2 implies that this skew-product is also minimal, so for each $s\in{\mathfrak{S}}^{\prime}$, the set $([w^{\prime}],s)$ reoccurs with a finite gap.
However, by a result of Durand [8] there are only finitely many derived spaces $X_{w^{\prime}}$ up to conjugacy, and thus only finitely many skew-products $F$. Hence, we can take the maximum over the gap bounds of all the finitely many (up to conjugacy) skew-products $F:X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}\to X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}$, and still have a finite constant.
Proof of Lemma 4.1. Let $Y=\bigcup_{x\in X_{w}}\{x\}\times s_{x}{\mathfrak{S}}_{x}$ be a transitive $F$-invariant subset of $X_{w}\times{\mathfrak{S}}$. From Proposition 3.1, we can take a word $w^{\prime}$ (longer than $w$ and associated with a cylinder set $U\subset X_{w}$) on which $s_{x}\equiv e$ and ${\mathfrak{S}}_{x}\equiv{\mathfrak{S}}^{\prime}$ is a constant subgroup of $\mathfrak{S}$}. Without loss of generality, we can assume that $U$ is given by a concatenation of return words of $w$ that itself is return word of some word $w^{\prime}$ in the original shift space $X$. Thus the return map to $U$ is the same as $(X_{w^{\prime}},\sigma)$ and the group-extension $F:X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}\to X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}$ is transitive.
Let $\mu_{w^{\prime}}$ be the unique shift-invariant measure on $X_{w^{\prime}}$ and $\nu^{\prime}$ counting measure on $\mathfrak{S}$$^{\prime}$. By Lemma 3.2, $\mu_{w^{\prime}}\times\nu^{\prime}|_{{\mathfrak{S}}^{\prime}}$ is ergodic, and then by Furstenberg’s Proposition 3.2, $F:X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}\to X_{w^{\prime}}\times{\mathfrak{S}}^{\prime}$ is also uniquely ergodic, and in particular minimal. Hence for each $s\in{\mathfrak{S}}^{\prime}$, the set $X_{w^{\prime}}\times\{s\}$ returns within a bounded number of steps. Let $L^{*}(X_{w},{\mathfrak{S}}^{\prime})<\infty$ be this bound maximized over the finitely many $s\in{\mathfrak{S}}^{\prime}$.
Durand [8] proved that there are only finitely many derived systems, so up to conjugacy, there are only finitely many systems $(X_{w^{\prime\prime}},{\mathfrak{S}}^{\prime})$, where $w^{\prime}$ is a prefix of $w^{\prime\prime}$. Thus, if we let $L^{*}$ be the maximal gap-length, maximized over all the finitely many (up to conjugacy) derived systems of $(X_{w^{\prime\prime}},{\mathfrak{S}}^{\prime})$, we have the uniform gap length claimed in Lemma 4.1. Therefore $L^{*}:=\max\{L^{*}(X_{w^{\prime}},{\mathfrak{S}}^{\prime}):w^{\prime}\in{\mathcal{L}}(X)\}$ satisfies the lemma for any $w^{\prime}$ associated to a derived system of $(X_{w},\sigma)$.
Now if $w^{\prime\prime}\in{\mathcal{L}}_{X}$ is any word, then by minimality of $(X,\sigma)$, there is always a word $w^{\prime}$ that we already treated and which has $w^{\prime\prime}$ as a subword. If we take $w^{\prime}$ of minimal length, then $|w^{\prime}|/|w^{\prime\prime}|$ is uniformly bounded as well, so the lemma follows. $\Box$
Proof of Theorem 1.1. Let $(X,\sigma)$ be a linearly recurrent subshift and $S$ a homeomorphic speedup. Let $v$ be some word for $S$; we refer to it as an $S$-pattern of the shift $(X,\sigma)$. Let $w$ be a sufficiently long word in $X$ such that center word $\hat{w}$ (see the construction of the group extension) of $w$ contains the $S$-pattern $v$. By linear recurrence, $w$ returns within $L|w|$ iterates, and by Lemma 4.1, the combination $(w,s)$ reoccurs within $L^{*}L|w|$ iterates. But the same $s\in{\mathfrak{S}}$ means that the $S$-orbits also return in the same order, so the $S$-orbit starting with $v$ sees $v$ again. Clearly $|w|\leq 2|\hat{w}|\leq 2p_{\max}|v|$, so $S$-pattern $v$ returns within $L^{*}L|w|\leq 2L^{*}Lp_{\max}|v|$ iterates of $S$. This gives the first implication of the theorem.
Now for the reverse implication, assume that $(X,S)$ is linearly recurrent with constant $L$, and let $v\in{\mathcal{L}}_{\sigma}(X)$ be any word. Recall from (1) that $p_{\max}=\max_{x\in X}p(x)<\infty$. Hence, there exist $N$ such that if $v^{\prime}=pvs$ is any extension of the word $v$ with an allowed prefix $p$ and suffix $s$ both of length $N$, then every $S$-orbits passing through $v^{\prime}$ passes through $v$ in a unique way. That is, let us write ergodic sums
[…]
then if $x\in[p.vs]$ (where the dot in the cylinder indicates the zero position), then $p_{k}(x)$ is constant on $[p.vs]$ for all $k\geq 0$ such that $p_{k}(x)\leq|v|$. Write the $S$-orbit of a chosen point $x\in[p.vs]$ as
$\dots w_{-n_{1}-1}w_{-n_{1}}\dots w_{-1}w_{0}w_{1}\dots w_{k}w_{k+1}\dots w_{k+n_{1}}w_{k+n_{1}+1}\dots$
where $w_{j}=x_{p_{j}(x)}x_{p_{j}(x)+1}\dots x_{p_{j+1}(x)-1}$ and $n_{1},n_{2}\leq N$ are chosen minimal such that $p_{-n_{1}}(x)\leq-N$ and $p_{k+n_{2}}(x)\geq|v|+N$. The word $\bar{w}=w_{-n_{1}}\dots w_{-1}w_{0}w_{1}\dots w_{k}w_{k+1}\dots w_{k+n_{1}}$ reappears in every $S$-orbit with gap $L|w|\leq Lp_{\max}(|v|+2N)\leq L^{\prime}|v|$ for some appropriate $L^{\prime}$ independent of $v$. But whenever $w$ appears in an $S$-orbit, $v$ appears as subword in the corresponding reconstructed $\sigma$-orbit. This shows that $(X,\sigma)$ is linearly recurrent with constant $L^{\prime}$. $\Box$
Acknowledgements HB was supported by the ANR-FWF Project I 6750-N. He also thanks Klaus Schmidt for useful discussions on non-abelian group extensions, and Jiao Changhua for communicating an issue with an earlier version of Lemma 3.2.